. A 0.010 mol sample of barium nitrate formed was heated until it had completely decomposed according to the following balanced equation:

2Ba (NO3)2(s) ˗˗˗˗˗˗˗˗˗˃˃ 2 BaO(s) + 4 NO2 (g) + O2(g)

a)     Use the ideal gas equation to calculate the total volume, cm3, of gaseous products obtained at 387 K and 1.12 × 105 Pa.

b)     The Kinetic Theory of gases is an attempt to explain the properties (behaviour) of ideal gases. What assumptions (at least three) does this theory make about gas molecules?


1
Expert's answer
2021-05-19T06:14:34-0400

a) Number of moles of the gas "=\frac{5}{2}*0.010=0.025 mol"

"V=\frac{nRT}{P}=\frac{0.025*0.0821*387}{1.12*10^5*9.86923*10^{-6}}=718.609307 cm^3"


b)The simplest kinetic model is based on the assumptions that:

(1) The gas is composed of a large number of identical molecules moving in random directions.

(2) The molecules undergo perfectly elastic collisions (no energy loss) with each other.

(3) The gas molecules are separated by distances that are large compared with their size.